OpenStudy (anonymous):

Find the graphing form, coordinates of the center & foci, the length of the transverse and conjugate axis, then graph the points and equation. 16x2 -81y2 + 324 y – 1620 = 0

OpenStudy (anonymous):

ok this one is not an ellipse is it? do you know what it is and how to tell?

OpenStudy (anonymous):

its asking for the length of the transverse and conjugate axis?

OpenStudy (anonymous):

lol maybe that is one way i had in mind that you have \(-81y^2\) not \(+81y^2\)

OpenStudy (anonymous):

we are going to do the same thing as before complete the square twice to make it look like \[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\]

OpenStudy (anonymous):

you think you can do that or you want me to walk you through it?

OpenStudy (anonymous):

walk me through it, i still need more practice

OpenStudy (anonymous):

ok there is only \(16x^2\) with no other \(x\) term so it will be easier factor out the \(-81\) from \[-81y^2+384y\] what do you get?

OpenStudy (anonymous):

\[16x^2-81(y^2+?)=1620\] what goes in the \(?\) spot? i.e. what is \(384\div(-81)\)

OpenStudy (anonymous):

-4.7

OpenStudy (anonymous):

whew i made a big mistake, i meant what is \(324\div (-81)\)

OpenStudy (anonymous):

i copied it down wrong sorry

OpenStudy (anonymous):

lol oh man, okay im confused now

OpenStudy (anonymous):

it was just a stupid typo on my part, don't let that throw you you wrote 324 and i copied it as 384

OpenStudy (anonymous):

so its -81y^2 - 324y?

OpenStudy (anonymous):

you have to factor out the \(-81\)

OpenStudy (anonymous):

\[16x^2-81(y^2-4y)=1620\]

OpenStudy (anonymous):

since \(-81\times -4=324\)

OpenStudy (anonymous):

did i lose you or is that okay now?

OpenStudy (anonymous):

i need you to get me back on track, what all i have written down is -81y^2 + 38y ,

OpenStudy (anonymous):

what do i do now?

OpenStudy (anonymous):

lets start from square one

OpenStudy (anonymous):

so delete that then?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

okay done

OpenStudy (anonymous):

lets start with \[16x^2 -81y^2 + 324y – 1620 = 0 \]add \(1620\) get \[16x^2-81y^2+324y=1620\] that part is easy

OpenStudy (anonymous):

now we are going to factor out the \(-81\) so we can complete the square

OpenStudy (anonymous):

\[16x^2-81(y^2-4y)=1620\] let me know when that is clear

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

ok good now to complete the square

OpenStudy (anonymous):

half of 4 is 2, so we write \[16x^2-81(y-2)^2=1620-?\]and we have to figure out what goes in the \(?\)

OpenStudy (anonymous):

\(y-2)^2=y^2-4y+4\) multiply that \(4\) by \(-81\) and you get \(324\) so you have to subtract \(324\) from the right \[16x^2-81(y-2)^2=1620-324\]

OpenStudy (anonymous):

\[16x^2-81(y-2)^2=1296\] then divide

OpenStudy (anonymous):

i have a feeling i might have lost you somewhere let me know if that is true, and where

OpenStudy (anonymous):

no you didnt i was just looking over what we're doing

OpenStudy (anonymous):

ok finally divide both sides by \(1296\) to put a 1 on the right

OpenStudy (anonymous):

this step always confuses me

OpenStudy (anonymous):

you do get that the right side will be 1 yes? when you divide by \(296\)?

OpenStudy (anonymous):

1296

OpenStudy (anonymous):

yes i got that part

OpenStudy (anonymous):

then these are usually cooked up to cancel nicely \[\frac{16}{1296}=\frac{1}{81}\]

OpenStudy (anonymous):

so the first term will be \[\frac{x^2}{81}\]

OpenStudy (anonymous):

and likewise \(\frac{81}{1296}=\frac{1}{16}\) so the next term will be \[-\frac{(y-2)^2}{16}\]

OpenStudy (anonymous):

then standard form is \[\frac{x^2}{81}-\frac{(y-2)^2}{16}=1\]

OpenStudy (anonymous):

now you can read off the center from your eyeballs again

OpenStudy (anonymous):

\[\frac{(x-h)^2}{a^2}-\frac{(y-k)^2}{b^2}=1\]\[\frac{x^2}{81}-\frac{(y-2)^2}{16}=1\]

OpenStudy (anonymous):

\((h,k)=(0,2)\)

OpenStudy (anonymous):

thats the center?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what else you need?

OpenStudy (anonymous):

foci and length of transverse and conjugate axis .-.

OpenStudy (anonymous):

ok again we need to know something does it look like the one on the left or on the right (lousy picture)